please kindly explain how -2x+20=0, becomes 0=0?

Found 2 solutions by Theo, Alan3354:
Answer by Theo(13342)

start with -2x + 20 = 0
add 2x to both sides of the equation to get:
20 = 2x
solve for x to get:
x = 10

here's where it happens.

to confirm your solution is correct, you replace x with the value that you found for it.

you get:

-2x + 20 = 0
replace x with 10 to get:
-2*10 + 20 = 0
simplify to get:
-20 + 20 = 0
combine like terms to get 0 = 0

since this is a true equation, your solution is good.

you confirm your solution is good by replacing the variable you solved for in the original equation and then determine if that equation holds true.

it holds true when the left side of the equation is equal to the right side of the equation.

0 = 0 is true because the left side of the equation actually is equal to the right side of the equation.

it just so happens that the right side of your original equation is equal to 0, so in order for the solution to be correct, the left side of the equation had to be equal to 0 as well once you solved for x.

for example:

suppose the equation was -2x + 20 = 36 instead.
you sould solve for x as follows:
add x to both sides of the equation and subtract 36 from both sides of the equation to get:
-16 = 2x
solve for x to get:
x = -8

now to confirm your solution is correct:
replace x with -8 in the original equation to get:
-2*-8 + 20 = 36
since -2 * -8 = 16, the equation becomes:
16 + 20 = 36
combine like terms to get:
36 = 36

since this equation is true, your solution is correct, because 36 actually is equal to 36 when you replace x with -8.

in the first case, 0 = 0 confirmed that solution to be good.
in the second case, 36 = 36 confirmed that solution to be good.

you always confirm your solution is correct, by replacing the variable solved for by the value that you solved it to be and then determine that the original equation is true when you do that.


Answer by Alan3354(69443)

please kindly explain how -2x+20=0, becomes 0=0?
========
It doesn't, there's no connection.
--------
You can solve -2x+20 = 0 for x --> x = 10
----------
0 = 0
1 = 1
75 = 75
etc.
There's no connection.